A rectangular parking lot is 50 ft longer than it is wide. Determine the dimensions of the parking lot if it measures 250 ft diagonally.
Answer by Prithwis(166)

Let the width be x ft.
The length is x+50 ft.
The length of the diagonal is sqrt [(x^2) + (x+50)^2] ft;
[(x^2) + (x+50)^2] = 250^2
=> x^2 + (x+50)^2 = 62500
=> x^2 + x^2 + 100x + 2500 = 62500
=> x^2 + 50x - 30000 = 0
=> x = [-50 + sqrt(2500+120000)]/2 because x cannot be negative
=> x = [-50 + sqrt(122500)]/2
=> x = (-50+350)/2 = 150;
The Dimension of the parking lot is 200 ft X 150 ft